# GUI‑AIMA: Aligning Intrinsic Multimodal Attention with a Context Anchor for GUI Grounding

### v1

@@ Abstract
Graphical user interface (GUI) grounding is a key function of computer-use agents, which maps natural-language instructions to actionable screen regions. Existing approaches based on Multimodal Large Language Models (MLLMs) typically formulate it as a text-based coordinate generation task, yet directly generating precise coordinates from visual inputs remains challenging and computationally intensive. An intuitive way to implement GUI grounding is to first select visual patches relevant to the instructions and then determine the precise click location within those patches. Based on the observations that general MLLMs have some native grounding capability, nested within their attentions, we propose GUI-AIMA, an attention-based and coordinate-free supervised fine-tuning framework for efficient GUI grounding. GUI-AIMA aligns the intrinsic multimodal attention of MLLMs with patch-wise grounding signals. These signals are calculated adaptively for diverse user instructions by multi-head aggregation on simplified query-visual attention matrices. Besides, its coordinate-free manner can easily integrate a plug-and-play zoom-in stage. GUI-AIMA-3B was trained with only 85k screenshots, demonstrating exceptional data efficiency and verifying that light training can trigger the native grounding capability of MLLMs. It achieves state-of-the-art performance among 3B models, attaining an average accuracy of 58.6% on ScreenSpot-Pro and 62.2% on OSWorld-G. Project page: https://github.com/sjz5202/GUI-AIMA.
@@ 1 Introduction
Graphical User Interface (GUI) (Hong et al., 2024; Cheng et al., 2024) agents have emerged as pivotal tools in automating interactions with digital devices, spanning from mobile applications (Rawles et al., 2023; 2024; Wang et al., 2024a; Ye et al., 2025a) to desktop software (Zhang et al., 2024; Qin et al., 2025; Deng et al., 2023; Xie et al., 2024; OpenAI, 2025; Zheng et al., 2024). GUI grounding plays an important role in mapping natural language instructions to specific elements on the screen, such as buttons, text fields, or icons (Yang et al., 2024). This process ensures that the agent’s actions are precise and contextually relevant, especially in environments with high-resolution displays and intricate layouts. However, the diversity of human instructions and GUI designs across platforms and applications (Li et al., 2025) makes GUI grounding a challenging task. Conventional methods rely on structured representations, such as HTML for web pages or accessibility trees (Zhang et al., 2025) for mobile applications. While these provide detailed information about the interface elements, they come with limited accessibility and excessive verbosity, which can lead to inefficiencies in processing. Moreover, structured data may omit essential visual cues such as layout and icons, which are critical for accurate grounding.
When humans use computers, they first determine the general area and then use visual feedback to interact with GUI elements (Ye et al., 2025b). Directly outputting coordinates from MLLMs might not be the most intuitive method. Instead, models should mimic human behavior:
@@
first selecting the relevant visual patch and then deciding the specific click position within that patch, i.e., the coordinates in the GUI. Previous studies, such as TAG (Xu et al., 2024a) and GUI-Actor (Wu et al., 2025), have explored the coordinate-free GUI grounding where corordinates are determined by selecting the most relevant visual patches in a screenshot instead of directly generating text tokens. However, GUI-Actor introduces extra modules into the MLLM, which requires an additional adaptation stage. Xu et al. (2024a) considers a vanilla but inaccurate way to identify and aggregate attention maps from all query text tokens, referred to as vanilla attention grounding.
In this work, we propose GUI-AIMA, an attention-based method for coordinate-free GUI grounding. We train the MLLM’s multi-head self-attention (MHSA) (Vaswani et al., 2017) with patch-wise grounding supervision using a novel attention head weighting mechanism. To simplify the impractical aggregation over all query tokens in vanilla attention grounding, we append a learnable <ANCHOR> token after the visual and text query tokens and supervise its attention over the visual tokens as a surrogate aggregator for all query tokens’ text-visual attentions. For precise multi-head aggregation, we weight each attention head by the magnitude of its query–visual correlation estimated from the MLLM’s visual-sink query tokens, rather than including visual-irrelevant query tokens as in vanilla attention grounding. The MLLM’s general visual-sink query tokens is identified by computing similarity between query and visual token embedding from layer-wise hidden states instead of relying on correlations within each individual attention head. These visual-sink query tokens guide the attention head weighting, prioritizing attention heads with strong interactions between visual-sink query and visual tokens, consistent with the general query-visual pattern revealed in hidden states, while down-weighting other heads. Together, compared with the vanilla attention grounding, GUI-AIMA achieves 4.5% improvements on ScreenSpot-pro. By operating directly on intrinsic attention without task-specific modules, GUI-AIMA enables data-efficient visual grounding training with only 85k screenshots. In addition, the center-encouraging and overlap-aware patch-wise labeling further boosts the performance of GUI-AIMA. Furthermore, offset grounding errors are prevalent on high-resolution screenshots due to dense visual content and detail loss from image compression in MLLMs. GUI-AIMA, with strong base grounding capacity for precise candidate-region localization, can combine the initially predicted region with an extra zoomed-in step without additional training to self-correct offset errors, as shown in Fig. 1.
@@ 2 Related Work
The core challenge for GUI agents (Wang et al., 2024b) is grounding: aligning user instructions with the correct actionable elements in the screenshot, due to the semantic inconsistency of layouts and UI elements across diverse GUI environments (Liu et al., 2025). In early attempts, along with the screenshots, extra structured inputs, such as HTML for U-Ground (Gou et al., 2024), or extractions from visual parsing modules, such as element extractions from OmniParser (Wan et al., 2024; Yu et al., 2025) and generated captions as contexts in Aria-UI (Yang et al., 2024), are fed into MLLMs as the supplementary inputs for the better interface understanding and more precise localization. Later works, such as AGUVIS (Xu et al., 2024b), SeeClick (Cheng et al., 2024) and OS-Atlas (Wu et al., 2024), explore the GUI grounding leveraging MLLMs (Bai et al., 2025; Chen et al., 2024b) with screenshot-only inputs, enabling the end-to-end localization with improved scalability across diverse GUI environments. The grounding approach in these methods is to generate coordinate-based click centers or bounding boxes as text, which is indirect alignment of visual grounding requiring additional efforts, such as scaled GUI corpora (Qin et al., 2025; Chai et al., 2024) and OCR pretraining (Hong et al., 2024) for connecting coordinates with UI elements. Recent endeavors manage to address this gap from the data-intensive manner of the coordinate-based GUI grounding methods via incorporating GUI-specific grounding reward signals into RL-training (Luo et al., 2025; Lu et al., 2025; Liu et al., 2025; Tang et al., 2025) and performing autonomous GUI exploration (Fan et al., 2025; Wu et al., 2025).
@@
Recent works have begun to replace traditional coordinate‑based grounding text with patch‑level attention map predictions, where patches belonging to the correct region receive higher weight. TAG (Xu et al., 2024a) first leveraged the multi-head self-attention between GUI instructions and visual tokens in MLLMs for tuning‑free GUI grounding, showcasing the intrinsic potential of MLLMs’ attention patterns for GUI tasks. However, the generalization of TAG for novel interfaces is restricted by the backbone and by the heuristics used for selecting text tokens and attention heads. SE-GUI (Yuan et al., 2025) retains a coordinate‑based prediction manner but employs self‑attention to iteratively filter noisy training samples during training. Aside these attention-based improvements, GUI-Actor (Wu et al., 2025) introduces an embedding‑based grounding head that derives attention between input visual patch embeddings and the final hidden state of a special <ACTOR> token. Compared with GUI-AIMA, GUI-Actor adds an additional module to connect cross-layer visual and <ACTOR> embeddings, requiring an extra warm-up training phase and resulting in much lower training efficiency.
@@ 3 Methods
The core idea of GUI-AIMA is utilizing the intrinsic query-visual pattern within attention maps of MLLMs. It performs supervised fine-tuning on the multi-head self-attention matrices across layers. Intuitively, as shown in Section 3.1, we can extract the attention vector between a text token and all visual patch tokens in each attention map of MLLM as the patch-wise attention vector, where each visual patch token’s attention weight can indicate its degree of membership in the grounding region. And the tuning-free aggregation of patch-wise attention vectors between query tokens has shown the potential for text localization (Xu et al., 2024a). Existing barriers for attention-based GUI grounding are how to aggregate patch-wise vectors for each query text tokens and each attention head. Besides, annotation gap between patch-wise and original coordinate labels further leads to inaccurate groundings. In GUI-AIMA, we propose a general attention-based GUI grounding framework that simplify the query token selection in Section 3.3 and refine the head-wise weighting mechanism for the patch-wise attention vector aggregation in Section 3.4. Furthermore, as detailed in Section 3.5, we extend GUI-AIMA with a two-step zoom-in inference that mitigates offset errors in grounding. Compared with previous embedding-based method, such as GUI-Actor (Wu et al., 2025), GUI-AIMA does not add extra grounding modules to MLLMs, eliminating the extra warm-up training phase.
@@ 3.1 Preliminary: Inherent GUI Grounding within Multi-head Self-Attention
where $w_{q_{i},(l,h)}$ is the aggregation weights for each query-visual attention vector with $w_{q_{i},(l,h)}\geq 0$ and $\sum_{l,h,i}w_{q_{i},(l,h)}=1$. We denote the aggregation in Eq. 1 as the Vanilla Attention Grounding shown as “vanilla” in Fig. 2. In the previous training-free attempt (Xu et al., 2024a) based on this manner, every selected text tokens are treated uniformly and attention heads are roughly selected. For the better adaptation of an attention-based GUI grounding method, we need a simplified and effective weighting strategy via $w_{q_{i},(l,h)}$.
@@ 3.2 Coordinate-free Patch-wise Labeling
The inherent grounding of MLLMs via the query-visual attention in Eq. 1 is patch-wise instead of the pixel manner in the original coordinate-based bounding boxes. In order to enable the patch-wise grounding supervision, we first reformat the ground truth bounding box $gt^{\mathrm{bbox}}=\left[x_{1},y_{1},x_{2},y_{2}\right]$ into patch-wise label vectors $\bm{p}\in\mathbb{R}^{|\mathcal{V}|}$, where $|\mathcal{V}|$ is the patch size of the interface $I$. Image patch $v_{i}$ is positive only if there is an overlap between $gt^{\mathrm{bbox}}$ and patch $v_{i}$. Considering the border patches which are partially overlapped with $gt^{\mathrm{bbox}}$ should be less weighted and the patches of center region should be highlighted, we weight each patch with $p_{v_{i}}$ according to the overlapping ratio of itself with $gt^{\mathrm{bbox}}$ and the distance from center of patch $v_{i}$ to the center of $gt^{\mathrm{bbox}}$ with Gaussian distribution following GUI-$\text{G}^{2}$ (Tang et al., 2025). In more details, we define the overlap between positive image patch $i$ and grounding box annotation $gt^{\mathrm{bbox}}$ using intersection over union ratio, denoted as $\textbf{IoU}(v_{i},gt^{\mathrm{bbox}})$. We take its center point $\bm{\mu}_{v_{i}}=(\text{center}^{v_{i}}_{x},\text{center}^{v_{i}}_{y})$ and the center point $\bm{\mu}_{gt}=(\text{center}^{gt}_{x},\text{center}^{gt}_{y})$ of $gt^{\mathrm{bbox}}$ for the following weighting scheme:
@@ 3.3 Efficient Grounding with a Visual Anchor Token
The supervision on $\hat{\mathbf{a}}$ in Eq. 1 requires a carefully calibrated balance across $\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}$ of each query token, where the weight of each query token $w_{q_{i},(l,h)}$ is difficult to determine. Moreover, direct supervision on the attention of query text tokens will impair the MLLM’s general capabilities (e.g., image captioning) and thus harm generalization (Nguyen et al., 2023).
@@
Figure 3: Details about how to compute the final patch-wise prediction based on the grounding vectors of <ANCHOR> token: GUI-AIMA first specifies visual-sink query tokens via computing hidden state similarities between query tokens and visual patches; Then it computes weights $\bm{w}$ of each attention head based on visual-sink query tokens in Eq. 8; Finally, GUI-AIMA aggregates the grounding vectors of <ANCHOR> token across layers and heads in Eq. 4.
@@ 3.4 Attention Head Weighting using Visual-sink Query Tokens
In Section 3.3, we introduce a special anchor token to avoid aggregating patch-wise attention vectors over all query text tokens. In this section, we continue to discuss how to determine $w_{l,h}$, the weights of different attention heads across different layers for the multi-head aggregation. Previous works (Li et al., 2023; Clark et al., 2019) has shown the functional diversity between attention heads across transformer blocks in LLMs. In GUI grounding, we want to identify attention heads, which shows active query-visual interactions. The attention head with large attention value between visual tokens $\mathcal{V}$ and text query tokens $\mathcal{Q}$ is more relevant, while others are not. As GUI grounding in the attention manner aims to capture the query-visual correspondence between the user instruction and interface, the above strategy focusing on query-visual interactions allows the attention supervision skew towards the attention heads with large query-visual correlations and affect the neutral attention heads less, thus improving the inherent attention grounding of MLLMs without impairing the pretrained capacity, as the adaptation complies with the original attention functionalities.
@@
Tokens with larger $c_{q_{i}}^{l}$ has strong visual correlations in layer $l$. The reason of measuring the query-visual affinity of $q_{i}$ in layer $l$ using hidden states $\mathbf{H}^{l}$ instead of attention, i.e. $c_{q_{i}}^{l}=\sum_{h\in[H]}\sum_{v_{i}\in\mathcal{V}}\,\mathbf{A}^{l,h}_{q_{i},v_{i}}$ is that the query-visual pattern discovered in hidden states $\mathbf{H}^{l}$ is not necessary statistically prevailing among each head’s self-attention matrix, as only a smaller subset of attention heads are “semantic heads” that key on semantic functionality and representation similarity reflected in $\mathbf{H}^{l}$ (Elhelo & Geva, 2024; Olsson et al., 2022; Voita et al., 2019) (an analysis is available in Appendix A). In comparison, the measurement in Eq. 7 is straightforward to reflect visual correlations of each text token in a global and general view. For grounding, we prioritize the “semantic” minority heads showing the similar pattern discovered from hidden states $\mathbf{H}$. Here, we denote the query tokens with $\operatorname{topK}$ large $c_{q_{i}}^{l}$ as the visual-sink query tokens $\mathcal{Q}_{s}^{l}$. The general query-visual correlation pattern of the MLLM captured by $\mathcal{Q}_{s}^{l}$ derived from hidden states should be consistent in the attention head heavily weighted for grounding for grounding, i.e., $\mathcal{Q}_{s}^{l}$ should also exhibit strong visual correlation in those heads. We measure the visual correlation on $\mathcal{Q}_{s}^{l}$ tokens in the individual attention head as the cumulative sum of attention values between $\mathcal{Q}_{s}^{l}$ and $\mathcal{V}$ in $\mathbf{A}^{l,h}$, as shown in the step 2 of Fig. 3, with further normalization:

### v2

# GUI-AIMA: Aligning Intrinsic Multimodal Attention with a Context Anchor for GUI Grounding

@@ Abstract
Graphical user interface (GUI) grounding is a key function of computer-use agents, which maps natural-language instructions to actionable screen regions. Existing approaches based on Multimodal Large Language Models (MLLMs) typically formulate it as a text-based coordinate generation task, yet directly generating precise coordinates from visual inputs remains challenging and computationally intensive. An intuitive way to implement GUI grounding is to first select visual patches relevant to the instructions and then determine the precise click location within those patches. Based on the observations that general MLLMs have some native grounding capability, nested within their attentions, we propose GUI-AIMA, an attention-based and coordinate-free supervised fine-tuning framework for efficient GUI grounding. GUI-AIMA aligns the intrinsic multimodal attention of MLLMs with patch-wise grounding signals. These signals are calculated adaptively for diverse user instructions by multi-head aggregation on simplified query-visual attention matrices. Besides, its coordinate-free manner can easily integrate a plug-and-play zoom-in stage. GUI-AIMA-3B was trained with only 85k screenshots, demonstrating exceptional data efficiency and verifying that light training can trigger the native grounding capability of MLLMs. It achieves state-of-the-art performance among 3B models, attaining an average accuracy of 59.6% on ScreenSpot-Pro, 63.8% on OSWorld-G and 91.5% on ScreenSpot-v2. Project page: https://github.com/sjz5202/GUI-AIMA.
@@ 1 Introduction
Graphical User Interface (GUI) (hong2024cogagent; cheng2024seeclick) agents have emerged as pivotal tools in automating interactions with digital devices, spanning from mobile applications (rawles2023androidinthewild; rawles2024androidworld; wang2024mobile; ye2025mobile) to desktop software (zhang2024ufo; qin2025ui; deng2023mind2web; xie2024osworld; cua2025; zheng2024gpt). GUI grounding plays an important role in mapping natural language instructions to specific elements on the screen, such as buttons, text fields, or icons (yang2024aria). This process ensures that the agent’s actions are precise and contextually relevant, especially in environments with high-resolution displays and intricate layouts. However, the diversity of human instructions and GUI designs across platforms and applications (li2025screenspot) makes GUI grounding a challenging task. Conventional methods rely on structured representations, such as HTML for web pages or accessibility trees (zhang2025appagent) for mobile applications. While these provide detailed information about the interface elements, they come with limited accessibility and excessive verbosity, which can lead to inefficiencies in processing. Moreover, structured data may omit essential visual cues such as layout and icons, which are critical for accurate grounding.
When humans use computers, they first determine the general area and then use visual feedback to interact with GUI elements (ye2025gui). Directly outputting coordinates from MLLMs might not be the most intuitive method. Instead, models should mimic human behavior:
@@
first selecting the relevant visual patch and then deciding the specific click position within that patch, i.e., the coordinates in the GUI. Previous studies, such as TAG (xu2024attentiondrivenguigroundingleveraging) and GUI-Actor (wu2025gui), have explored the coordinate-free GUI grounding where corordinates are determined by selecting the most relevant visual patches in a screenshot instead of directly generating text tokens. However, GUI-Actor introduces extra modules into the MLLM, which requires an additional adaptation stage. xu2024attentiondrivenguigroundingleveraging considers a vanilla but inaccurate way to identify and aggregate attention maps from all query text tokens, referred to as vanilla attention grounding.
In this work, we propose GUI-AIMA, an attention-based method for coordinate-free GUI grounding. We train the MLLM’s multi-head self-attention (MHSA) (vaswani2017attention) with patch-wise grounding supervision using a novel attention head weighting mechanism. To simplify the impractical aggregation over all query tokens in vanilla attention grounding, we append a learnable <ANCHOR> token after the visual and text query tokens and supervise its attention over the visual tokens as a surrogate aggregator for all query tokens’ text-visual attentions. For precise multi-head aggregation, we weight each attention head by the magnitude of its query–visual correlation estimated from the MLLM’s visual-sink query tokens, rather than including visual-irrelevant query tokens as in vanilla attention grounding. The MLLM’s general visual-sink query tokens is identified by computing similarity between query and visual token embedding from layer-wise hidden states instead of relying on correlations within each individual attention head. These visual-sink query tokens guide the attention head weighting, prioritizing attention heads with strong interactions between visual-sink query and visual tokens, consistent with the general query-visual pattern revealed in hidden states, while down-weighting other heads. Together, compared with the vanilla attention grounding, GUI-AIMA achieves 4.5% improvements on ScreenSpot-pro. By operating directly on intrinsic attention without task-specific modules, GUI-AIMA enables data-efficient visual grounding training with only 85k screenshots. In addition, the center-encouraging and overlap-aware patch-wise labeling further boosts the performance of GUI-AIMA. Furthermore, offset grounding errors are prevalent on high-resolution screenshots due to dense visual content and detail loss from image compression in MLLMs. GUI-AIMA, with strong base grounding capacity for precise candidate-region localization, can combine the initially predicted region with an extra zoomed-in step without additional training to self-correct offset errors, as shown in Fig. 1.
@@ 2 Related Work
The core challenge for GUI agents (wang2024gui) is grounding: aligning user instructions with the correct actionable elements in the screenshot, due to the semantic inconsistency of layouts and UI elements across diverse GUI environments (liu2025infiguig1advancingguigrounding). In early attempts, along with the screenshots, extra structured inputs, such as HTML for U-Ground (gou2024navigating), or extractions from visual parsing modules, such as element extractions from OmniParser (Wan_2024_CVPR; yu2025omniparserv2structuredpointsofthoughtunified) and generated captions as contexts in Aria-UI (yang2024aria), are fed into MLLMs as the supplementary inputs for the better interface understanding and more precise localization. Later works, such as AGUVIS (xu2024aguvis), SeeClick (cheng2024seeclick) and OS-Atlas (wu2024atlas), explore the GUI grounding leveraging MLLMs (bai2025qwen2; chen2024expanding) with screenshot-only inputs, enabling the end-to-end localization with improved scalability across diverse GUI environments. The grounding approach in these methods is to generate coordinate-based click centers or bounding boxes as text, which is indirect alignment of visual grounding requiring additional efforts, such as scaled GUI corpora (qin2025ui; chai2024amex) and OCR pretraining (hong2024cogagent) for connecting coordinates with UI elements. Recent endeavors manage to address this gap from the data-intensive manner of the coordinate-based GUI grounding methods via incorporating GUI-specific grounding reward signals into RL-training (luo2025gui; lu2025ui; liu2025infiguig1advancingguigrounding; tang2025gui) and performing autonomous GUI exploration (fan2025gui; wu2025gui).
@@
Recent works have begun to replace traditional coordinate‑based grounding text with patch‑level attention map predictions, where patches belonging to the correct region receive higher weight. TAG (xu2024attentiondrivenguigroundingleveraging) first leveraged the multi-head self-attention between GUI instructions and visual tokens in MLLMs for tuning‑free GUI grounding, showcasing the intrinsic potential of MLLMs’ attention patterns for GUI tasks. However, the generalization of TAG for novel interfaces is restricted by the backbone and by the heuristics used for selecting text tokens and attention heads. SE-GUI (yuan2025enhancing) retains a coordinate‑based prediction manner but employs self‑attention to iteratively filter noisy training samples during training. Aside these attention-based improvements, GUI-Actor (wu2025gui) introduces an embedding‑based grounding head that derives attention between input visual patch embeddings and the final hidden state of a special <ACTOR> token. Compared with GUI-AIMA, GUI-Actor adds an additional module to connect cross-layer visual and <ACTOR> embeddings, requiring an extra warm-up training phase and resulting in much lower training efficiency.
@@ 3 Methods
The core idea of GUI-AIMA is utilizing the intrinsic query-visual pattern within attention maps of MLLMs. It performs supervised fine-tuning on the multi-head self-attention matrices across layers. Intuitively, as shown in Section 3.1, we can extract the attention vector between a text token and all visual patch tokens in each attention map of MLLM as the patch-wise attention vector, where each visual patch token’s attention weight can indicate its degree of membership in the grounding region. And the tuning-free aggregation of patch-wise attention vectors between query tokens has shown the potential for text localization (xu2024attentiondrivenguigroundingleveraging). Existing barriers for attention-based GUI grounding are how to aggregate patch-wise vectors for each query text tokens and each attention head. Besides, annotation gap between patch-wise and original coordinate labels further leads to inaccurate groundings. In GUI-AIMA, we propose a general attention-based GUI grounding framework that simplify the query token selection in Section 3.3 and refine the head-wise weighting mechanism for the patch-wise attention vector aggregation in Section 3.4. Furthermore, as detailed in Section 3.5, we extend GUI-AIMA with a two-step zoom-in inference that mitigates offset errors in grounding. Compared with previous embedding-based method, such as GUI-Actor (wu2025gui), GUI-AIMA does not add extra grounding modules to MLLMs, eliminating the extra warm-up training phase.
@@ 3.1 Preliminary: Inherent GUI Grounding within Multi-head Self-Attention
where $w_{q_{i},(l,h)}$ is the aggregation weights for each query-visual attention vector with $w_{q_{i},(l,h)}\geq 0$ and $\sum_{l,h,i}w_{q_{i},(l,h)}=1$. We denote the aggregation in Eq. 1 as the Vanilla Attention Grounding shown as “vanilla” in Fig. 2. In the previous training-free attempt (xu2024attentiondrivenguigroundingleveraging) based on this manner, every selected text tokens are treated uniformly and attention heads are roughly selected. For the better adaptation of an attention-based GUI grounding method, we need a simplified and effective weighting strategy via $w_{q_{i},(l,h)}$.
@@ 3.2 Coordinate-free Patch-wise Labeling
The inherent grounding of MLLMs via the query-visual attention in Eq. 1 is patch-wise instead of the pixel manner in the original coordinate-based bounding boxes. In order to enable the patch-wise grounding supervision, we first reformat the ground truth bounding box $gt^{\mathrm{bbox}}=\left[x_{1},y_{1},x_{2},y_{2}\right]$ into patch-wise label vectors $\bm{p}\in\mathbb{R}^{|\mathcal{V}|}$, where $|\mathcal{V}|$ is the patch size of the interface $I$. Image patch $v_{i}$ is positive only if there is an overlap between $gt^{\mathrm{bbox}}$ and patch $v_{i}$. Considering the border patches which are partially overlapped with $gt^{\mathrm{bbox}}$ should be less weighted and the patches of center region should be highlighted, we weight each patch with $p_{v_{i}}$ according to the overlapping ratio of itself with $gt^{\mathrm{bbox}}$ and the distance from center of patch $v_{i}$ to the center of $gt^{\mathrm{bbox}}$ with Gaussian distribution following GUI-$\text{G}^{2}$ (tang2025gui). In more details, we define the overlap between positive image patch $i$ and grounding box annotation $gt^{\mathrm{bbox}}$ using intersection over union ratio, denoted as $\textbf{IoU}(v_{i},gt^{\mathrm{bbox}})$. We take its center point $\bm{\mu}_{v_{i}}=(\text{center}^{v_{i}}_{x},\text{center}^{v_{i}}_{y})$ and the center point $\bm{\mu}_{gt}=(\text{center}^{gt}_{x},\text{center}^{gt}_{y})$ of $gt^{\mathrm{bbox}}$ for the following weighting scheme:
@@ 3.3 Efficient Grounding with a Visual Anchor Token
The supervision on $\hat{\mathbf{a}}$ in Eq. 1 requires a carefully calibrated balance across $\mathbf{A}^{l,h}_{q_{i},\mathcal{V}}$ of each query token, where the weight of each query token $w_{q_{i},(l,h)}$ is difficult to determine. Moreover, direct supervision on the attention of query text tokens will impair the MLLM’s general capabilities (e.g., image captioning) and thus harm generalization (nguyen2023regularization).
@@
Figure 3: Details about how to compute the final patch-wise prediction based on the grounding vectors of the <ANCHOR> token: GUI-AIMA first specifies visual-sink query tokens via computing hidden state similarities between query tokens and visual patches; Then it computes weights $\bm{w}$ of each attention head based on visual-sink query tokens in Eq. 8; Finally, GUI-AIMA aggregates the grounding vectors of <ANCHOR> token across layers and heads in Eq. 4.
@@ 3.4 Attention Head Weighting using Visual-sink Query Tokens
In Section 3.3, we introduce a special anchor token to avoid aggregating patch-wise attention vectors over all query text tokens. In this section, we continue to discuss how to determine $w_{l,h}$, the weights of different attention heads across different layers for the multi-head aggregation. Previous works (li2023inference; clark2019does) has shown the functional diversity between attention heads across transformer blocks in LLMs. In GUI grounding, we want to identify attention heads, which shows active query-visual interactions. The attention head with large attention value between visual tokens $\mathcal{V}$ and text query tokens $\mathcal{Q}$ is more relevant, while others are not. As GUI grounding in the attention manner aims to capture the query-visual correspondence between the user instruction and interface, the above strategy focusing on query-visual interactions allows the attention supervision skew towards the attention heads with large query-visual correlations and affect the neutral attention heads less, thus improving the inherent attention grounding of MLLMs without impairing the pretrained capacity, as the adaptation complies with the original attention functionalities.
@@
Tokens with larger $c_{q_{i}}^{l}$ has strong visual correlations in layer $l$. The reason of measuring the query-visual affinity of $q_{i}$ in layer $l$ using hidden states $\mathbf{H}^{l}$ instead of attention, i.e. $c_{q_{i}}^{l}=\sum_{h\in[H]}\sum_{v_{i}\in\mathcal{V}}\,\mathbf{A}^{l,h}_{q_{i},v_{i}}$ is that the query-visual pattern discovered in hidden states $\mathbf{H}^{l}$ is not necessary statistically prevailing among each head’s self-attention matrix, as only a smaller subset of attention heads are “semantic heads” that key on semantic functionality and representation similarity reflected in $\mathbf{H}^{l}$ (elhelo2024inferring; olsson2022context; voita2019analyzing) (an analysis is available in Appendix A). In comparison, the measurement in Eq. 7 is straightforward to reflect visual correlations of each text token in a global and general view. For grounding, we prioritize the “semantic” minority heads showing the similar pattern discovered from hidden states $\mathbf{H}$. Here, we denote the query tokens with $\operatorname{topK}$ large $c_{q_{i}}^{l}$ as the visual-sink query tokens $\mathcal{Q}_{s}^{l}$. The general query-visual correlation pattern of the MLLM captured by $\mathcal{Q}_{s}^{l}$ derived from hidden states should be consistent in the attention head heavily weighted for grounding for grounding, i.e., $\mathcal{Q}_{s}^{l}$ should also exhibit strong visual correlation in those heads. We measure the visual correlation on $\mathcal{Q}_{s}^{l}$ tokens in the individual attention head as the cumulative sum of attention values between $\mathcal{Q}_{s}^{l}$ and $\mathcal{V}$ in $\mathbf{A}^{l,h}$, as shown in the step 2 of Fig. 3, with further normalization: